Solve the equation 4^2x+1=2^3x+6.
Possible solutions are:
a. -5
b. 2
c. 4
d. 5

Answer by Nate(3500)

4^(2x + 1) = 2^(3x + 6)
2^(4x + 2) = 2^(3x + 6)
4x + 2 = 3x + 6
x = 4
(c)